2.3. Let G be a nonempty set closed under an associative product, which in addition satisfies:

(a) There exists an e ∈ G such that ae = a for all a ∈ G.

(b) Given a ∈ G, there exists an element y(a) ∈ G such that ay(a) = e.

Prove that G must be a group under this product.


1
Expert's answer
2022-06-20T15:26:04-0400

Given "a\in G". Since right inverse exists, there exists "y(a)\in G" such that "ay(a)=e". Then, "y(a)=y(a)e=y(a)(ay(a))=(y(a)a)y(a)". Also, there exists "t\in G" such that "y(a)t=e". This implies that "(y(a)a)y(a)t=e" then "(y(a)a)=e". Hence "y(a)a=e". So every right inverse is also a left inverse.

Now for any "a\in G" we have "ea=(ay(a))a=a(y(a)a)=ae=a" as "e" is a right identity. Hence "e" is left identity.